OpenStudy (anonymous):

Solve 4 log12 2 + log12 x = log12 96. x = 88 x = 80 x = 12 x = 6

OpenStudy (anonymous):

@IMStuck

OpenStudy (australopithecus):

Whenever you see log, \[\log(x) = \log_{10}(x)\] Now for the laws you need to solve this problem: \[\log(a) + \log(b) = \log(ab)\] \[\log(a^b) = b*\log(a)\]

OpenStudy (australopithecus):

Oh sorry forgot one, \[a^{\log_a(x)} = x\] For example, \[12^{\log_{12}(x)} = x\] \[5^{\log_{5}(x)} = x\]

OpenStudy (imstuck):

Ok, when you end up with \[\log _{12}\frac{ 16x }{ 96 }=0\]that translates to \[12^{0}=\frac{ 16x }{ 96 }\]

OpenStudy (imstuck):

That is because of this rule:

OpenStudy (imstuck):

\[\log _{a}x=y-->x=a ^{y}\]

OpenStudy (imstuck):

That's the hardest part of this one.

OpenStudy (imstuck):

Anything to the 0 power = 1, so we have \[1=\frac{ 16x }{ 96 }\]

OpenStudy (imstuck):

Now multiply both sides by 96 to get 96=16x

OpenStudy (imstuck):

Now divide by 16 to get x = 6, your last choice up there.

OpenStudy (imstuck):

This one was a stinky one. : P

OpenStudy (anonymous):

@IMStuck your a lifesaver

OpenStudy (imstuck):

TY for the medal, as well!